OpenStudy (anonymous):

A firecracker shoots up from a hill 155 feet high with an initial speed of 100 feet per second. Using the formula H(t) = -16t2 + vt + s, determine how long it will take the firecracker to hit the ground.

OpenStudy (eujc21):



OpenStudy (phi):

you have an equation that tells you the height of the firecracker. \[ H(t) = -16t^2 + vt + s\] You have s in this formula. We have to put in a number. What is the height of the firecracker at t=0 (right before it zooms off)? sub in 0 wherever you see t in the formula: H(0)= -16*0*0+v*0+s which simplifies to H(0)= s so s is the height of the firecracker at the very beginning. What is its height? from a hill 155 feet high I would say 155 ft (the top of the hill). We also know from the problem that v= initial speed of 100 feet per second. so the equation is \[ H(t)= -16t^2+100t+155 \] how long it will take the firecracker to hit the ground? set H(t)=0 (0 is the height when you are on the ground) \[ 0= -16t^2+100t+155 \] can you solve a quadratic equation for t?